# Towards More Robust Retrieval-Augmented Generation: Evaluating RAG Under Adversarial Poisoning Attacks

### v1

@@ D.2 Patterns of Expressing Uncertainty
GPT-3.5 Adversarial 4.00 0.00$\scriptstyle(\downarrow 4.00)$ 4.00 0.00$\scriptstyle(\downarrow 4.00)$ 4.00 0.00$\scriptstyle(\downarrow 4.00)$
Untouched 6.00 4.00$\scriptstyle(\downarrow 2.00)$ 9.00 5.00$\scriptstyle(\downarrow 4.00)$ 3.00 2.00$\scriptstyle(\downarrow 1.00)$
GPT-4 Adversarial 10.00 1.00$\scriptstyle(\downarrow 9.00)$ 14.00 0.00$\scriptstyle(\downarrow 14.00)$ 9.00 0.00$\scriptstyle(\downarrow 9.00)$
Untouched 33.00 8.00$\scriptstyle(\downarrow 25.00)$ 3.00 1.00$\scriptstyle(\downarrow 2.00)$ 17.00 5.00$\scriptstyle(\downarrow 12.00)$
GPT-4o Adversarial 13.00 1.00$\scriptstyle(\downarrow 12.00)$ 6.00 0.00$\scriptstyle(\downarrow 6.00)$ 5.00 0.00$\scriptstyle(\downarrow 5.00)$
Untouched 16.00 16.00(0.00)0.00$\scriptstyle$(0.00)( 0.00 ) 1.00 3.00$\scriptstyle(\uparrow 2.00)$ 7.00 3.00$\scriptstyle(\downarrow 4.00)$
Llama-8b Adversarial 6.00 0.00$\scriptstyle(\downarrow 6.00)$ 13.00 0.00$\scriptstyle(\downarrow 13.00)$ 5.00 0.00$\scriptstyle(\downarrow 5.00)$
Untouched 23.00 9.00$\scriptstyle(\downarrow 14.00)$ 24.00 $\scriptstyle(\downarrow 17.00)$ 6.00 $\scriptstyle(\downarrow 3.00)$
Llama-70b Adversarial 7.00 0.00$\scriptstyle(\downarrow 7.00)$ 8.00 0.00$\scriptstyle(\downarrow 8.00)$ 8.00 0.00$\scriptstyle(\downarrow 8.00)$
Untouched 20.00 11.00$\scriptstyle(\downarrow 9.00)$ 21.00 4.00$\scriptstyle(\downarrow 17.00)$ 9.00 4.00$\scriptstyle(\downarrow 5.00)$
Claude-3.5 Adversarial 64.00 12.00$\scriptstyle(\downarrow 52.00)$ 35.00 1.00$\scriptstyle(\downarrow 34.00)$ 25.00 0.00$\scriptstyle(\downarrow 25.00)$
Untouched 56.00 43.00$\scriptstyle(\downarrow 13.00)$ 37.00 2.00$\scriptstyle(\downarrow 35.00)$ 34.00 5.00$\scriptstyle(\downarrow 29.00)$
@@
D.3 Effect of retrieval pollution rate
@@ D.4 Effect of increasing $k$
ionCounteract Adversarial Passages In Figure 13 and 14, we illustrate the heatmap of F1 score with different number of adversarial passage and guiding passage, the x-axis is the number of adversarial passage and y axis is the number of guiding passage. The first row is the result for neutral prompting and the second row is the result for skeptical prompting. In Figure 16 and Figure 15, we have F1 score changing with the number of adversarial context with 0, 2, and 4 guiding context to counteract the effect of adversarial context.
D.5 F1 Scores for Faithful Prompting and Guiding Contexts
@@ 20 Adv 96 100 90 94 93
D.6 Combined retrieval and generation result for HotpotQA dataset

### v2

@@ E.2 Patterns of Expressing Uncertainty
GPT-3.5 Adversarial 4.00 0.00(4.00$\scriptstyle$($\downarrow$ 4.00)( ↓ 4.00 ) 4.00 0.00(4.00$\scriptstyle$($\downarrow$ 4.00)( ↓ 4.00 ) 4.00 0.00(4.00$\scriptstyle$($\downarrow$ 4.00)( ↓ 4.00 )
Untouched 6.00 4.00(2.00$\scriptstyle$($\downarrow$ 2.00)( ↓ 2.00 ) 9.00 5.00(4.00$\scriptstyle$($\downarrow$ 4.00)( ↓ 4.00 ) 3.00 2.00(1.00$\scriptstyle$($\downarrow$ 1.00)( ↓ 1.00 )
GPT-4 Adversarial 10.00 1.00(9.00$\scriptstyle$($\downarrow$ 9.00)( ↓ 9.00 ) 14.00 0.00(14.00$\scriptstyle$($\downarrow$ 14.00)( ↓ 14.00 ) 9.00 0.00(9.00$\scriptstyle$($\downarrow$ 9.00)( ↓ 9.00 )
Untouched 33.00 8.00(25.00$\scriptstyle$($\downarrow$ 25.00)( ↓ 25.00 ) 3.00 1.00(2.00$\scriptstyle$($\downarrow$ 2.00)( ↓ 2.00 ) 17.00 5.00(12.00$\scriptstyle$($\downarrow$ 12.00)( ↓ 12.00 )
GPT-4o Adversarial 13.00 1.00(12.00$\scriptstyle$($\downarrow$ 12.00)( ↓ 12.00 ) 6.00 0.00(6.00$\scriptstyle$($\downarrow$ 6.00)( ↓ 6.00 ) 5.00 0.00(5.00$\scriptstyle$($\downarrow$ 5.00)( ↓ 5.00 )
Untouched 16.00 16.00(0.00$\scriptstyle$(0.00)( 0.00 ) 1.00 3.00(2.00$\scriptstyle$($\uparrow$ 2.00)( ↑ 2.00 ) 7.00 3.00(4.00$\scriptstyle$($\downarrow$ 4.00)( ↓ 4.00 )
Llama-8b Adversarial 6.00 0.00(6.00$\scriptstyle$($\downarrow$ 6.00)( ↓ 6.00 ) 13.00 0.00(13.00$\scriptstyle$($\downarrow$ 13.00)( ↓ 13.00 ) 5.00 0.00(5.00$\scriptstyle$($\downarrow$ 5.00)( ↓ 5.00 )
Untouched 23.00 9.00(14.00$\scriptstyle$($\downarrow$ 14.00)( ↓ 14.00 ) 24.00 7.00(17.00$\scriptstyle$($\downarrow$ 17.00)( ↓ 17.00 ) 6.00 $\scriptstyle(\downarrow 3.00)$
Llama-70b Adversarial 7.00 0.00(7.00$\scriptstyle$($\downarrow$ 7.00)( ↓ 7.00 ) 8.00 0.00(8.00$\scriptstyle$($\downarrow$ 8.00)( ↓ 8.00 ) 8.00 0.00(8.00$\scriptstyle$($\downarrow$ 8.00)( ↓ 8.00 )
Untouched 20.00 11.00(9.00$\scriptstyle$($\downarrow$ 9.00)( ↓ 9.00 ) 21.00 4.00(17.00$\scriptstyle$($\downarrow$ 17.00)( ↓ 17.00 ) 9.00 4.00(5.00$\scriptstyle$($\downarrow$ 5.00)( ↓ 5.00 )
Claude-3.5 Adversarial 64.00 12.00(52.00$\scriptstyle$($\downarrow$ 52.00)( ↓ 52.00 ) 35.00 1.00(34.00$\scriptstyle$($\downarrow$ 34.00)( ↓ 34.00 ) 25.00 0.00(25.00$\scriptstyle$($\downarrow$ 25.00)( ↓ 25.00 )
Untouched 56.00 43.00(13.00$\scriptstyle$($\downarrow$ 13.00)( ↓ 13.00 ) 37.00 2.00(35.00$\scriptstyle$($\downarrow$ 35.00)( ↓ 35.00 ) 34.00 5.00(29.00$\scriptstyle$($\downarrow$ 29.00)( ↓ 29.00 )
@@
E.3 Effect of retrieval pollution rate
@@ E.4 Effect of increasing $k$
E.5 Counteract Adversarial Passages
In Figure 13 and 14, we illustrate the heatmap of F1 score with different number of adversarial passage and guiding passage, the x-axis is the number of adversarial passage and y axis is the number of guiding passage. The first row is the result for neutral prompting and the second row is the result for skeptical prompting. In Figure 16 and Figure 15, we have F1 score changing with the number of adversarial context with 0, 2, and 4 guiding context to counteract the effect of adversarial context.
E.6 F1 Scores for Faithful Prompting and Guiding Contexts
@@ 20 Adv 96 100 90 94 93
E.7 Combined retrieval and generation result for HotpotQA dataset